OpenStudy (anonymous):

The given measurements may or may not determine a triangle. If not, then state that no triangle is formed. If a triangle is formed, then use the Law of Sines to solve the triangle, if it is possible, or state that the Law of Sines cannot be used. B = 111°, c = 8, b = 12

OpenStudy (anonymous):

options: C = 38.5°, A = 30.5°, a ≈ 6.5 No triangle is formed. C = 30.5°, A = 38.5°, a ≈ 6.5 The triangle cannot be solved with the Law of Sines.

OpenStudy (anonymous):

@satellite73 can you help me with this?

OpenStudy (anonymous):

yeah sure

OpenStudy (anonymous):

\[\frac{\sin(111)}{12}=\frac{\sin(C)}{8}\] from the law of sines

OpenStudy (anonymous):

this means \[\sin(C)=\frac{8\sin (111)}{12}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

if \(\frac{8\sin(111)}{12}\) is bigger than one, there is no solution, because sine is never bigger than one

OpenStudy (anonymous):

it isn't.

OpenStudy (anonymous):

but really what we want is not \[\sin(C)\] but rather \(C \) itself

OpenStudy (anonymous):

and so \(C\) is the arcsine of all that mess

OpenStudy (anonymous):

i.e. \[C=\sin^{-1}\left(\frac{8\sin(111)}{12}\right)\]

OpenStudy (anonymous):

38.49

OpenStudy (anonymous):

=C

OpenStudy (anonymous):

if you say so

OpenStudy (anonymous):

lol what do you mean by that

OpenStudy (anonymous):

i mean i didn't do it, but you are right

OpenStudy (anonymous):

oh ok thanks :D

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

So is the 3rd answer choice correct?

OpenStudy (anonymous):

C = 38.5°, A = 30.5°, a ≈ 6.5